OpenStudy (anonymous):

Marty has a piggy bank that contains $2.25. The piggy bank contains only quarters and dimes. There are twice as many dimes as quarters in the bank. How many dimes are in the piggy bank?

OpenStudy (anonymous):

10d + 25q = 225 d = 2q 10(2q) +25q = 225 20q + 25q = 225 45q = 225 q = 5 d = 10 Answer: 5 quartters ($1.25) + 10 dimes($1.00) are in the piggy bank.

OpenStudy (anonymous):

Thanks:)